# SensorChat: Answering Qualitative and Quantitative Questions during Long-Term Multimodal Sensor Interactions

### v1

@@ 4.2. Contrastive Sensor-Text Pretraining for Partial Contexts
Figure 5. Visualization of the contrastive sensor-text pretraining in SensorChat.
We introduce a new pretraining loss to align sensor embeddings $\mathbf{z}^{s}_{t}$ and partial text embeddings $\mathbf{z}^{w}_{t}$. Different from CLIP (Radford et al., 2021), our loss function enables alignment between sensor data and all partial phrases, defined as follows:
@@
4.3. Three Stages in SensorChat
In this section, we explain the detailed designs in each stage of SensorChat, including question decomposition, sensor data query and answer assembly. All stages need to work accurately, robustly and collaboratively to achieve natural and precise answer generation in the end.
4.3.1. Question Decomposition
The question decomposition stage processes the user question and identifies specific query triggers for the sensor data query stage, such as the context of interest, date and time ranges, and the summarization function, as illustrated in the leftmost box in Fig. 4. The primary challenge in designing this stage lies in generating accurate decompositions for arbitrary user inputs, including various questions types as discussed in Sec. 3. We choose to rely on LLMs for this stage due to their outstanding capability in handling natural language inputs. However, LLMs are not without limitations as they can produce erroneous outputs or hallucinations (Huang et al., 2023), especially in our case where a dedicated dataset for similar decomposition tasks is unavailable.
To address this challenge, we propose a few-shot learning approach to prompt pre-trained LLMs, enhanced with in-context learning (Alayrac et al., 2022; Shao et al., 2023) and chain of thought techniques (Chu et al., 2024; Lu et al., 2022). An example prompt is illustrated in Fig. 6(a), where LLM is instructed to mark different arguments with distinct symbols, such as ”¡¡” and ”¿¿” for function names. This helps in accurately extracting various arguments from the LLM output. In SensorChat, we accommodate a variety of real-life scenarios by extracting contexts (such as activities), dates, times of day, and summarization functions during the query process, as shown in Fig. 6(a). SensorChat supports multiple extracted phrases to handle complex questions such as ”How long did I work at school on Monday and Tuesday?” SensorChat is designed to be flexible, allowing for future expansion with additional decomposition terms.
We next explain more details on the in-context learning and chain-of-thought techniques used in SensorChat, both for ensuring high-quality question decompositions.
In-Context Learning (ICL) integrates a few examples directly into the prompt during inference, enabling LLMs to adapt effectively to specific tasks without requiring fine-tuning (Alayrac et al., 2022; Shao et al., 2023). Building on this insight, we design a library of general solution templates covering diverse QA scenarios and incorporate them as in-context examples to improve decomposition accuracy (Fig. 6(a)). We observe that decomposition solutions for the same question type often share similarities. For instance, ”how long” questions typically map to the CalculateDuration function. To utilize this question-specific property, in SensorChat, we first classify the question type using a BERT model (Devlin et al., 2018). Depending on the question type, we dynamically select templates that best match the question category, a strategy shown to enhance ICL performance (Fu et al., 2023). Fig. 6(b) illustrates an example of a solution template for a time query. Including an explanation in the solution template is particularly helpful for complex reasoning tasks.
Chain-of-Thought (CoT) Prompting explicitly requests LLMs to generate their reasoning process, which improves accuracy in zero- or few-shot logical reasoning (Chu et al., 2024; Lu et al., 2022). In SensorChat, we include “please generate step-by-step explanations” in prompts to encourage CoT, as shown in Fig. 6(a). CoT and detailed explanations in solution templates improve the quality of logical reasoning, especially for questions requiring multi-step thoughts. For example, applying CoT to the question “What did I do right after waking up on Wednesday?” helps produce a step-by-step thinking process, such as: “To find the activity right after waking up on Wednesday, we can query the start and end times of all activities on Wednesday, therefore we can use the function ¡¡DetectFirstTime¿¿ and ¡¡DetectLastTime¿¿ with ((all activities)) on [[Wednesday]]”.
(a) The prompt design for the question decomposition stage includes ICL and CoT.
(b) Example solution template for the time query question category.
(c) Prompt design in the answer assembly stage for LLM finetuning.
Figure 6. Prompts design in SensorChat to better leverage the power of LLMs.
4.3.2. Sensor Data Query
The sensor query stage is the core module of SensorChat, responsible for searching the sensor database and extracting relevant information, which directly impacts the accuracy of SensorChat’s answers. Existing sensor-based QA systems did not employ an explicit search module, restricting to low-dimensional sensor features (Englhardt et al., 2024; Kim et al., 2024; Yang et al., 2024b) or fixed windows of sensor signals (Xing et al., 2021; Moon et al., 2023b, a; Chen et al., 2024; Arakawa et al., 2024). In contrast, SensorChat introduces this explicit query stage to handle long-duration, fine-grained timeseries sensor signals and provide accurate information to queries, which is not possible with prior QA designs.
The sensor and label encoders are pretrained offline as explained in Sec. 4.2. With the pretrained sensor encoder, the full-history raw sensor signals are converted into an embedding database beforehand. During online user interactions, as shown in the middle part of Fig. 4, the sensor query stage only needs to encode the context (e.g., activity) generated in the question decomposition stage using the text encoder (④ in Fig. 4) , performs a similarity search among the sensor database given the text (⑤), and summarizes the retrieved sensor embeddings into textual information for the final answer assembly (⑥). The design of the sensor query stage faces two primary challenges: (1) ensuring accurate and efficient searches within the sensor embedding database, and (2) constructing adequate output context based on users’ queries. SensorChat addresses these challenges through an efficient similarity-based embedding search mechanism and a set of carefully designed summarization functions, as detailed below.
@@ 4.3.2. Sensor Data Query
Similarity-based Embedding Search Our goal is to accurately identify all sensor samples relevant to the query context. For example, if the query context is “exercise”, SensorChat aims to retrieve all sensor samples associated with activities similar to exercise. If multiple text are generated during question decomposition, SensorChat performs an embedding search for each query text individually. The pretraining in Sec. 4.2 ensures that the sensor embedding space is well aligned with the text embedding space even using partial context. Therefore, the similarity comparison given any pair of sensor and text embeddings can be achieved by training a similarity function $f$ in the embedding space:
@@
The output of $f$ is a scalar value between 0 and 1, representing the similarity between the sensor and text samples. With $f$, the similarity search in the embedding space is significantly more efficient than searching directly through raw, high-dimensional time-series data. The efficiency can be further improved by narrowing the search scope based on the date and time-of-day arguments identified during question decomposition.
Summarization Function Design We design a set of summarization functions to ”summarize” the queried sensor samples and generate contextual information for the answer assembly stage. The specific summarization function to be used is determined by SensorChat during question decomposition. For example, a question of ”how long” should be directed to the CalculateDuration function, while a question of ”what did I do” should be handled by the DetectActivity function. Each summarization function uses a unique template to return text information. The numerical value in the returned text is determined by the embedding search results. For instance, when querying CalculateDuration with the activity “cooking”, date “Sunday”, and time “morning”, the text output can be: ”You spent $\gamma$ minutes cooking on Sunday morning”, where $\gamma$ is calculated as follows:
@@
Here $T_{SundayMorning}$ represents all timestamps within Sunday morning. The notation $[Cond]$ gives $1$, when the inner condition Cond is met; otherwise 00. $h$ is a predetermined threshold.
In SensorChat, we carefully design a set of summarization functions to account for diverse scenarios in real life, including time quries, activity quries, counting, etc. The details of those functions are explained in Appendix A.
4.3.3. Answer Assembly
As shown in the right box of Fig. 4, the final stage of answer assembly integrates question and sensor information to generate the final answer. State-of-the-art methods (Xing et al., 2021; Zhang et al., 2023; Moon et al., 2023b) rely on “black-box” fusion of natural language and sensory data, often leading to ineffective fusion and inaccurate answers (see Sec. 3). In contrast, SensorChat summarizes query results to text and directly fuses them with the question in the prompt, as illustrated in Fig. 6(c). Our intuition is that, in contrast to processing and fusing with other modalities, LLMs are the most professional in dealing with text. SensorChat is capable of answering both qualitative and quantitative questions by combining the original question and the extracted fine-grained activity information from long-duration, high-dimensional sensors. At this answer assembly stage, we finetune a LLM such as LLaMA (Zhang et al., 2023) to adapt the model to the desired answer style. Fine-tuning is chosen over few-shot learning as it delivers better performance with the presence of high-quality datasets like SensorQA (Reichman et al., 2025) (see Sec.6.6). We use Low Rank Adaptation (LoRA) (Hu et al., 2021) due to its parameter efficiency and comparable performance to full fine-tuning.
@@ 5. System Implementation
We implement SensorChat on real-world systems. We envision SensorChat as a personal assistant that provides accurate and timely answers to user questions, as outlined in our problem statement (Sec. 3.1). Fig. 7 visualizes the general system pipeline of SensorChat. We employ smartphones and smartwatches to collect multimodal sensor data from users in daily lives (① in Fig. 7(a)). Implemented based on the ExtraSensory App (Vaizman et al., 2018a), the mobile devices automatically gathers data for 20 seconds every minute, including 40Hz IMU signals, 13 MFCC audio features from a 22kHz microphone, and other phone state information including compass, GPS location, Wi-Fi status, light intensity, battery level, etc. The details can be found in the ExtraSensory App manual (Vaizman et al., 2018a). These sensor data are then transmitted to a system running SensorChat (② in Fig. 7(a)). We offer two variants of SensorChat, designed for a cloud server and an edge environment respectively. Their detailed implementations and trade-offs are discussed below. Finally, users can interact with SensorChat directly through a chatting interface using natural language, shown as ③ in Fig. 7(a).
SensorChatC and SensorChatE We offer two system variants of SensorChat, as shown in Fig. 7(b) and 7(c).
@@
SensorChatC and SensorChatE accommodate two typical use scenarios. SensorChatC is expected to deliver superior QA performance with the full-precision LLMs in the cloud, at the cost of intensive resource consumption. Additionally, SensorChatC requires the users to transmit the full sensor history to the cloud server. On the other hand, SensorChatE runs entirely on a local edge platform belonging to the user, eliminating the need to transmit user data to the cloud and thus preserving user privacy. However, its QA and latency performance degrade compared to SensorChatC.
@@
Figure 7. System implementation details of SensorChat.
Implementation Details The algorithm part of SensorChat is implemented with Python and PyTorch (Paszke et al., 2019). For the offline sensor encoder, SensorChat uses a Transformer architecture (Vaswani, 2017) with 6 encoder layers, 8 attention heads, and a feedforward network size of 2048 for time series data. For the low-dimensional phone status data, SensorChat employs a fully connected layer as the encoder. The fusion layer is also implemented as a fully connected layer. The label encoder is initialized from the pretrained CLIP ViT-B/32 label encoder (Radford et al., 2021). Both the sensor and label embedding spaces share a dimension size of 512. Offline pretraining is performed with the partial-context loss proposed in Sec. 4.2 using a temperature scalar of $\tau=0.1$. We use the Adam optimizer with a learning rate of $51e-$ 5 over 100 epochs.
For question decomposition, we design two solution templates for each question category. Limiting the number of templates to two helps balance prompt length and performance. For online data queries, we initialize the classifier $f$ as a multilayer perceptron with one hidden layer and a ReLU activation function. The hidden layer size is set to 512 and the query threshold is configured to $h=0.5$. For finetuning the LLM in answer assembly, we apply low rank adaptation (LoRA) finetuning (Hu et al., 2021) on an A100 GPU, using a batch size of 8, a learning rate of 0.0002 and 10 epochs. The LoRA rank $r$ is 16, the LoRA scaling factor alpha is 16, and the LoRA dropout is 0.1. During answer generation, SensorChat uses a max sequence length of 1024 and a geberating temperature of $0.2$.
@@ 6. Evaluation on State-of-the-Art Dataset
In this section, we thoroughly evaluate SensorChat on the state-of-the-art dataset focusing on quantitative questions. We further validate SensorChat in a real world study with open-ended, qualitative questions in Sec. 7.
@@ 6.1. Dataset and Metrics
In this evaluation section, we focus mainly on the SensorQA dataset (Reichman et al., 2025) and quantitative questions. A detailed introduction to SensorQA (Reichman et al., 2025) is provided in Sec. 3. To the best of our knowledge, SensorQA is the first and only available benchmarking dataset for QA interactions that use long-term timeseries sensor data and reflect practical user interests. While we focus on SensorQA (Reichman et al., 2025) in this section, we emphasize that the motivation and design of SensorChat are broadly applicable and can be extended to other practical sensing applications. To ensure the best alignment between the QA pairs and sensor information, we conduct offline encoders pretraining on the ExtraSensory multimodal sensor dataset (Vaizman et al., 2017), which servers as the sensor data source for SensorQA. During pretraining, all sensor samples are aligned by a time window of 20 seconds.
@@
  • Short answers are the 1-2 key words extracted from the full answers by GPT-3.5-Turbo (Ye et al., 2023), offered with the original SensorQA dataset (Reichman et al., 2025). We use the exact match accuracy on the short answers to evaluate the precision of generated answers, as detailed in Sec. 3.
@@
We create the multiple-choice version in addition to the full answers and short answers provided in the original SensorQA dataset (Reichman et al., 2025), to further assess the model’s ability in distinguishing similar facts based on sensor data. We use different dataset variants to assess various aspects of the models. The full answers dataset evaluates overall language quality, while the short answers and multiple-choice evaluations focus on the model’s ability to learn underlying facts rather than relying solely on patterns in language token generation. Further discussion of the evaluation metrics can be found in Sec. 8.
@@ 6.2. State-of-the-Art Baselines
We compare SensorChat against several state-of-the-art baselines that leverage different modalities combinations, including text, vision+text, and sensor+text data. These comparisons highlight the effectiveness of integrating multiple modalities to address the sensor-based QA tasks. We consider both closed-source and open-source baselines for a comprehensive analysis.
We first evaluate the state-of-the-art closed-source generative models using various modalities:
  • GPT-3.5-Turbo (Ye et al., 2023) and GPT-4 (Achiam et al., 2023) are text-only baselines taking only the questions as inputs.
  • GPT-4-Turbo (Achiam et al., 2023) and GPT-4o (Achiam et al., 2023) are vision+text baselines taking the activity graphs and the questions as inputs. For all vision+text baselines, we feed the activity graphs in SensorQA (Reichman et al., 2025), similar to Fig. 2, along with the questions into the model.
  • IMU2CLIP+GPT-4222https://github.com/facebookresearch/imu2clip (Moon et al., 2023b) is the state-of-the-art sensor+text GPT baseline as explained in Sec. 3.
For these closed-source generative models, we use few-shot learning (FSL). Specifically, we incorporate a set of QA examples from SensorQA (Reichman et al., 2025) into the prompt for each question input. We adopt $10$ samples per question based on a grid search of $\{2,5,10,15\}$.
We further conduct comprehensive evaluation with the state-of-the-art open-source models using various modalities:
  • T5 (Raffel et al., 2020) and LLaMA [^3] https://www.llama.com/ (Touvron et al., 2023a) are popular text-only language models.
  • LLaMA-Adapter [^4] https://github.com/OpenGVLab/LLaMA-Adapter (Zhang et al., 2023) is a recent vision+text framework offering a lightweight method for fine-tuning instruction-following and multimodal LLaMA models. It integrates vision inputs (i.e., activity graphs from SensorQA (Reichman et al., 2025)) with LLMs using a transformer adapter. We utilize the latest LLaMA-Adapter V2 model.
  • LLaVA-1.5555https://llava-vl.github.io/ (Liu et al., 2024a) represents state-of-the-art vision+text model. LLaVA connects pre-trained CLIP ViT-L/14 visual encoder (Radford et al., 2021) and large language model Vicuna (Chiang et al., 2023), using a projection matrix. LLaVA-1.5 (Liu et al., 2024a) achieves state-of-the-art performance on 11 benchmarks through simple modifications to the original LLaVA and the use of more extensive public datasets for finetuning.
  • DeepSQA [^6] https://github.com/nesl/DeepSQA (Xing et al., 2021) trains a CNN-LSTM model with compositional attention to fuse sensor+text modalities and predict from a fixed and limited set of candidate answers given questions and IMU signals. We adapt their implementation to use the full-history timeseries data as input, to align with SensorChat’s setup.
  • OneLLM [^7] https://github.com/csuhan/OneLLM (Han et al., 2023) is a state-of-the-art multimodal LLM framework that processes sensor+text modalities using a universal pretrained CLIP encoder and a mixture of projection experts for modality alignment. We adapt their implementation and feed the full-history timeseries data from the IMU tokenizer.
@@
All baselines based on LLaMA (that is, LLaMA, LLaMA-Adapter, and OneLLM) use LLaMA2-7B (Touvron et al., 2023a). For T5 and DeepSQA, we train the models directly on the SensorQA dataset (Reichman et al., 2025). For the LLM baselines, we apply LoRA fine-tuning (FT) (Hu et al., 2021) using the samples from SensorQA (Reichman et al., 2025). For all methods that require training, we randomly select 80% of the QA pairs in SensorQA (Reichman et al., 2025) as training samples and reserve the remaining 20% for testing. We explore alternative splitting schemes in Sec. 6.8 to demonstrate SensorChat’s generalizability to unseen users. All baselines adopt the same hyperparameters as those specified in their official codebases. We did not compare with the latest works of Sensor2Text (Chen et al., 2024), PrISM-Q&A (Arakawa et al., 2024), and DrHouse (Yang et al., 2024b) due to limited access to their open-source code and models.
(a) Example of a time query question on a single-day duration.
(b) Example of a counting question on a multi-day duration.
Figure 8. Qualitative results of SensorChat in comparison to state-of-the-art methods.
6.3. Qualitative Performance
We illustrate a qualitative comparison of SensorChat with the top-performing baselines in Fig. 8. Specifically, Fig. 8 (a) focuses on time-related queries and (b) focuses on counting questions, which are the most challenging for SOTA methods given the long-duration time series sensor inputs (see Sec. 3). Key phrases in the answers are highlighted in the green if they closely match the true answer, and in the red if they do not. The two presented examples are very challenging for state-of-the-art baselines. SensorChat, on the other hand, consistently produces more accurate answers which can be attributed to its novel three-stage pipeline.
In Fig. 8 (b), counting the total days spent at home requires long-term reasoning where LLaMA-Adapter (Zhang et al., 2023) and DeepSQA (Xing et al., 2021) usually fall short. SensorChat decomposes this question and queries the duration of ”at home” on ”each day”, then leaving the counting task to the answering stage. In a nutshell, SensorChat relies on question decomposition and sensor data queries to extract relevant key sensor information, while the answer assembly stage handles reasoning and produces the final answer. This collaboration across the three stages allows SensorChat to effectively manage a wide range of tasks, particularly those requiring multi-step reasoning and quantitative analyzes, which highlights SensorChat’s advancements over existing works.

### v2

@@ 4.3.1. Contrastive Sensor-Text Pretraining for Partial Contexts
[…]
We introduce a new pretraining loss to align sensor embeddings $\mathbf{z}^{s}_{t}$ and partial text embeddings $\mathbf{z}^{w}_{t}$. Different from CLIP (Radford et al., 2021), our loss function enables alignment between sensor data and all partial phrases:
@@
$q^{i},a^{i}$ Question & answer pair $c^{i}$ Extracted context from question decomposition
@@
4.3.2. Online Sensor Data Query
SensorChat further introduces a query mechanism to extract relevant sensor embeddings and convert them into textual summarizations based on the pretrained sensor and text encoders. The active query process of SensorChat differs significantly from the passive journaling or summarization modules (Xu et al., 2024b; Ouyang and Srivastava, 2024), as SensorChat dynamically selects the query functions and query ranges based on the questions. The performance of sensor data query directly impacts the accuracy of SensorChat’s answers. As shown in the middle part of Fig. 3, after SensorChat encodes the context $c_{i}$ using the text encoder (④), SensorChat performs a similarity search among the sensor embedding given the text embedding (⑤), and summarizes the retrieved sensor embeddings into textual information using custom query functions (⑥). We next detail the similarity-based embedding search and query function design.
@@ 4.3.2. Online Sensor Data Query
Similarity-based Embedding Search Our goal is to accurately identify all sensor samples relevant to the query context. For example, if the query context $c_{i}$ is “exercise”, SensorChat aims to retrieve all sensor samples associated with activities similar to exercise. If multiple text are generated during question decomposition, SensorChat performs an embedding search for each query text individually. The pretraining in Sec. 4.3.1 ensures that the sensor embedding space is well aligned with the text embedding space even using partial context. Therefore, the similarity comparison given any pair of sensor and text embeddings can be achieved by training a similarity function $f$ in the embedding space:
@@
The output of $f$ is a scalar value between 0 and 1, representing the similarity between the sensor and text samples. With $f$, the similarity search in the embedding space is significantly more efficient than searching directly through raw, high-frequency time-series data. The efficiency can be further improved by narrowing the search scope based on the date and time-of-day arguments identified during question decomposition.
Query Function Design SensorChat uses a set of query functions to filter the relevant sensor samples and generate textual information for the answer assembly stage. The specific query function to be used is determined during question decomposition. For example, a question of ”how long” should be directed to the CalculateDuration function, while a question of ”what did I do” should be handled by the DetectActivity function. Each query function uses a unique template to return text information. The numerical value in the returned text is determined by the embedding search results. For instance, when querying CalculateDuration with the activity “cooking”, date “Sunday”, and time “morning”, the textual output can be: ”You spent $\gamma$ minutes cooking on Sunday morning”, where $\gamma$ is calculated as follows:
@@
Here $T_{SundayMorning}$ represents all timestamps within Sunday morning. The notation $[Cond]$ gives $1$, when the inner condition Cond is met; otherwise 00. $h$ is a predetermined threshold. In SensorChat, we carefully design a set of query functions to account for diverse scenarios in real life, including time quries, activity quries, counting, etc. The details of those functions are explained in Appendix B.
4.4. Answer Assembly
As shown in the right box of Fig. 3, the final stage of answer assembly integrates question and sensor information to generate the final answer. State-of-the-art methods (Xing et al., 2021; Zhang et al., 2023; Moon et al., 2023b) rely on training a neural network to fuse natural language and sensor data, similar to a “black-box”. These methods often lead to ineffective fusion and inaccurate answers when dealing with long-duration, high-frequency sensor data (see Sec. 3.2). In contrast, SensorChat directly fuses the filtered senor context from the query stage with the original question. Our intuition is that, in contrast to processing and fusing with other modalities, LLMs are the most professional in dealing with text. SensorChat is capable of answering both qualitative and quantitative questions by combining the original question and the extracted fine-grained activity information from long-duration, high-frequency sensors. At this answer assembly stage, we finetune a LLM such as LLaMA (Zhang et al., 2023) to adapt the model to the desired answer style. Fine-tuning is chosen over few-shot learning as it delivers better performance with the presence of high-quality datasets like SensorQA (Reichman et al., 2025), which is verified in Appendix F. We use Low Rank Adaptation (LoRA) (Hu et al., 2021) due to its parameter efficiency and comparable performance to full fine-tuning. The prompt is as follows: “Answer the question based on the context below. Context: {Context} Question: {Question} Response:”
@@ 5. System Implementation
We implement SensorChat on real-world systems. We envision SensorChat as a personal assistant that provides accurate and timely answers to user questions, as outlined in our problem statement (Sec. 3.1). Fig. 6 visualizes the general system pipeline of SensorChat. We employ smartphones and smartwatches to collect multimodal sensor data from users in daily lives (① in Fig. 6(a)). Implemented based on the ExtraSensory App (Vaizman et al., 2018), the mobile devices automatically gathers data for 20 seconds every minute, including 40Hz IMU signals, 13 MFCC audio features from a 22kHz microphone, and other phone state information including compass, GPS location, Wi-Fi status, light intensity, battery level, etc. The details can be found in the ExtraSensory App manual (Vaizman et al., 2018). These sensor data are then transmitted to a system running SensorChat (② in Fig. 6(a)). We offer two variants of SensorChat, designed for a cloud server and an edge environment respectively. Their detailed implementations and trade-offs are discussed below. Finally, users can interact with SensorChat directly through a chatting interface using natural language, shown as ③ in Fig. 6(a).
SensorChatC and SensorChatE We offer two system variants of SensorChat, as shown in Fig. 6(b) and 6(c).
@@
SensorChatC and SensorChatE accommodate two typical use scenarios. SensorChatC is expected to deliver superior QA performance with the full-precision LLMs in the cloud, at the cost of intensive resource consumption. Additionally, SensorChatC requires the users to transmit the full sensor history to the cloud server. On the other hand, SensorChatE runs entirely on a local edge platform belonging to the user, eliminating the need to transmit user data to the cloud and thus preserving user privacy. However, its QA and latency performance degrade compared to SensorChatC. Further implementation details can be found in Appendix C.
@@
Figure 6. System implementation of SensorChat.
@@ 6. Evaluation on State-of-the-Art Dataset
In this section, we thoroughly evaluate SensorChat on the state-of-the-art dataset focusing on quantitative questions. Due to space limitations, we defer the ablation study, sensitivity analysis, and generalization study to Appendix FI. We further validate SensorChat in a real world study with both quantitative and qualitative questions in Sec. 7.
@@ 6.1. Dataset and Metrics
In this evaluation section, we focus mainly on the SensorQA dataset (Reichman et al., 2025) and quantitative questions. To the best of our knowledge, SensorQA is the first and only available benchmarking dataset for QA interactions that use long-term timeseries sensor data and reflect practical user interests. While we focus on SensorQA (Reichman et al., 2025) in this section, we emphasize that SensorChat is broadly applicable and can be extended to other real-world sensing applications. To ensure the best alignment between the QA pairs and sensor information, we conduct offline encoders pretraining on the ExtraSensory multimodal sensor dataset (Vaizman et al., 2017), which servers as the sensor data source for SensorQA. During pretraining, all sensor samples are aligned by a time window of 20 seconds.
@@
  • Short answers are the 1-2 key words extracted from the full answers by GPT-3.5-Turbo (Ye et al., 2023), offered with the original SensorQA dataset (Reichman et al., 2025). We use the exact match accuracy on the short answers to evaluate the precision of generated answers, as detailed in Sec. 3.2.
@@
We create the multiple-choice version in addition to the full answers and short answers provided in the original SensorQA dataset (Reichman et al., 2025), to further assess the model’s ability in distinguishing similar facts based on sensor data.
@@ 6.2. State-of-the-Art Baselines
We compare SensorChat against the state-of-the-art baselines that leverage different modalities combinations, including text, vision+text, and sensor+text data. For the text-only models, we employ the popular GPT-4 (Achiam et al., 2023), T5 (Raffel et al., 2020) and LLaMA (Touvron et al., 2023a) to obtain a baseline performance without considering the sensor data. For the vision+text models, we adopt GPT-4-Turbo (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), LLaMA-Adapter (Zhang et al., 2023) and LLaVA-1.5 (Liu et al., 2024a). We feed the activity graphs in SensorQA (Reichman et al., 2025) (similar to Fig. 1(c)) into these models along with the questions as input. We also compare against the latest sensor+text models including IMU2CLIP+GPT-4o (Moon et al., 2023b), LLMSense (Ouyang and Srivastava, 2024), DeepSQA (Xing et al., 2021) and OneLLM (Han et al., 2023). To ensure a fair comparison, we adapt these baselines to process the same long-duration, high-frequency data by modifying the data format to match their original input modality. Further adaptation details are provided in Appendix D.
@@
For the closed-source generative baselines based on GPT, we use few-shot learning (FSL). Specifically, we incorporate a set of QA examples from SensorQA (Reichman et al., 2025) into the prompt for each question input. We adopt $10$ samples per question based on a grid search of $\{2,5,10,15\}$. For T5 and DeepSQA, we train the models directly on the SensorQA dataset (Reichman et al., 2025). For open-source LLM baselines, we apply LoRA fine-tuning (FT) (Hu et al., 2021) using the samples from SensorQA (Reichman et al., 2025). We randomly select 80% of the QA pairs in SensorQA (Reichman et al., 2025) as training samples and reserve the remaining 20% for testing. We explore alternative splitting schemes in Appendix I to demonstrate SensorChat’s generalizability to unseen users. All baselines based on LLaMA (that is, LLaMA, LLaMA-Adapter, and OneLLM) use LLaMA2-7B (Touvron et al., 2023a). All baselines adopt the same hyperparameters as those specified in their official codebases.
Figure 7. Qualitative results of SensorChat in comparison to state-of-the-art methods on SensorQA (Reichman et al., 2025). Key phrases in the answers are highlighted in the green if they match the true answer, and in the red if they do not. SensorChat is the only method among the presented baselines that correctly interprets the question and generates an accurate answer.
6.3. QA Performance
We start with presenting an example of SensorChat compared to top-performing baselines in Fig. 7. Additional QA examples are presented in Appendix E. Answering the question in Fig. 7 involves two steps: counting meal frequency per day and identifying the day with the fewest meals. All three baselines fail in this challenging task. DeepSQA (Xing et al., 2021) and LLaMA-Adapter (Zhang et al., 2023) struggle with long-duration, high-frequency data, as they were originally designed for short sensor inputs under one minute. In our setting, these models fail to learn effective fusion between long-term sensor data and textual questions, leading to wrong answers. IMU2CLIP+GPT-4o (Moon et al., 2023b) works by first mapping raw sensor signals into a long narrative log and then passing the log to GPT-4o for reasoning. Although IMU2CLIP+GPT-4o is able to accurately associate sensor signals with text, it is constrained by GPT-4o’s token limit, causing key portions of the log to be truncated and leading to incomplete or uninformative answers. In contrast, SensorChat succeeds by decomposing the query in the first stage, extracting daily meal frequencies in the second stage, and leaving the comparison to the answer stage. The collaboration across the three stages allows SensorChat to effectively handle quantitative queries that require precise numerical reasoning across multi-day span, which highlights SensorChat’s advancements over existing works.